# Inverse Reinforcement Learning by Estimating Expertise of Demonstrators

### v1

@@ 4.3 Extension to Generalized IRL
where $\psi$ is a convex reward regularizer that smoothly penalizes differences between occupancy measures to prevent overfitting (Ho & Ermon, 2016). A naive solution to this nested min-max objective involves an outer loop learning the reward, and an inner loop executing maximum entropy RL with this reward to find the optimal policy. As before, the optimal policy under reward $r$ takes the form of the soft Bellman policy defined in Equation 3, where $r(s,a)$ replaces the linear reward $\theta^{\top}f(s,a)$. The description of IRLEED in the generalized setting follows directly.
@@
where in practice, the unknown functions $r$ and $\epsilon=\{\epsilon_{i}\}_{i=1}^{N}$ can be parameterized by neural networks. Note that under a constant reward regularizer $\psi$, if we express the rewards as $r_{i}(s,a)=(\theta+\epsilon_{i})^{\top}f(s,a)$, then the gradients of the above objective are identical to the ones described in Lemma 4.2.
In summary, we have shown how any learning algorithm that solves the generalized IRL objective defined in Equation 15 can be utilized alongside our framework by re-paremetrizing the reward and policy accordingly. Our experiment results in Section 5 demonstrate that IRLEED performs well in this generalized setting, and can even be trained offline when paired with Inverse soft-Q learning (IQ), an approach that avoids the iterative process defined by learning a single Q-function, implicitly representing both reward and policy (Garg et al., 2021).
5 Experiments
In this section, we empirically evaluate how IRLEED performs when learning from suboptimal demonstrations, using experiments in both online and offline IL settings, with simulated and human-generated data. In total, we performed three sets of experiments, comparing IRLEED to maximum entropy IRL (Ziebart et al., 2010), Inverse soft-Q learning (IQ) (Garg et al., 2021), and ILEED (Beliaev et al., 2022). We go over the results of each experiment below, briefly detailing the corresponding setup for each. For further implementation details refer to Appendix A.
5.1 Gridworld
For our first experiment, we used a custom gridworld environment, simulating suboptimal demonstrations with varying levels of precision and accuracy. Using this environment, we compared the performance of IRLEED against maximum entropy IRL.
Setup The reward in this setting was linearly parameterized as $r(s)=\theta^{\top}f(s)$, where $f(s)$ is simply a one hot encoding for each gridworld state. To implement IRL, we utilized stochastic value iteration on the reward estimate $\theta$ to infer the policy $\pi$, and Monte Carlo simulations of the MDP to estimate the feature expectations required to optimize $\theta$, repeating both steps until convergence. We used the same processes to implement IRLEED according to the procedure outlined in Section 4, updating $\epsilon$ and $\beta$ only after $\theta$ converged, and repeating this for two iterations.
For each level of precision and accuracy, we randomly generated $5$ suboptimal policies according to Equation 6, where $\beta_{i}$ was sampled from a uniform distribution starting at 00, with mean equal to the precision level, and $\epsilon_{i}$ was sampled from a multivariate normal distribution $\mathcal{N}([0]^{k},\mathbf{I}_{k}/\lambda^{2})$, with $\lambda$ equal to the accuracy level. We tested a total of $121$ dataset settings, collecting $40$ trajectories from each policy to compose our dataset, and comparing the performance of IRLEED and IRL over $100$ seeds for each setting.
Results We first compare how IRLEED and IRL recover the reward $r(s,a)$ when provided with suboptimal demonstrations. We visualize the averaged result for one dataset setting in Figure 2, which shows the true reward, the dataset quality, and the recovered rewards under both methods. We can see that the provided demonstrations are misaligned with the ground truth reward: the state visitation frequency for the top left corner is higher due to demonstrator suboptimalities. As expected, the feature matching constraint of IRL absorbs this suboptimality. Although the reward recovered by IRL contains information about the true reward, it is now incorrectly biasing the top left corner. On the other hand, IRLEED is able to remove this bias, providing us with a better estimate of the ground truth reward.
To see the effect of dataset quality, we use Figure 3 to show the relative improvement of the policy recovered by IRLEED over maximum entropy IRL for each dataset setting. We can see from the left plot that IRLEED provides more improvement over IRL as we decrease accuracy $\epsilon$ (by decreasing $\lambda$), or increase precision $\beta$. This signifies the importance of modeling the precision $\epsilon$, or reward bias, in addition to the accuracy $\beta$, or action variance, when learning from suboptimal demonstrations. This point is further highlighted in the right plot, which shows that although both IRLEED and IRL suffer when demonstrator accuracy is diminished, IRLEED can tolerate lower demonstrator accuracy compared to IRL. In total, IRLEED provided a $30.3\%$ improvement over IRL for the $121$ settings tested.
Figure 2: We visualize the reward recovered by IRLEED and IRL when trained using suboptimal demonstrations. Top left shows the true reward, where the three yellow corners are terminal states. Top right shows the normalized state visitation frequency over the entire dataset. Bottom left and right show the normalized rewards recovered by IRLEED and IRL respectively.
Figure 3: We compare the performance of the policies recovered by IRLEED and IRL. The left plot shows the relative performance of IRLEED over IRL under varying dataset settings, where the top right corner corresponds to expert data. The right plot shows the performance of both policies as we increase the average accuracy $\epsilon$ of demonstrators in the dataset, corresponding to the data outlined in red in the left plot.
5.2 Simulated Control Tasks
For our second experiment, we used two control tasks, Cartpole, and Lunar Lander, simulating suboptimal demonstrations with varying levels of expertise and accuracy. Using these two environments, we compared the performance of IRLEED, IQ, and ILEED in both the online and offline settings.
Setup To implement IQ, and setup the training and evaluation procedures, we utilized the codebase provided by the authors. We implemented ILEED alongside this codebase, using the training and evaluation procedures. For IRLEED, we built on top of the aforementioned IQ implementation by adding a learnable parameter $\beta_{i}$, and an additional critic network $\epsilon_{i}$, for each demonstrator $i$. Since IQ learns a reward and policy directly by using a single Q function as a critic, we used our additional critic network $\epsilon_{i}$ to directly add bias to the state action function instead of the reward.
For this experiment, we used $4$ dataset settings corresponding to the precision and accuracy levels: (high, high), (high, low), (low, high), and (low, low). For each setting, we randomly generated $3$ suboptimal policies according to $\pi_{i}(a|s)\propto\exp(\beta_{i}(Q(s,a)+\alpha\epsilon_{i}(s,a)))$, where $Q(s,a)$ is a pretrained critic network, $\epsilon_{i}(s,a)$ is a randomly initialized critic network, and $\alpha$ is a scaling factor which is set to zero for the high accuracy setting. For the low precision setting, $\beta_{i}$ was sampled from a uniform distribution, whereas for the high precision setting, the stochasticity of the policy was removed by following the maximum state action value. We ran this experiment in both the online and offline setting, comparing the performance of IRLEED, IQ, and ILEED, using $30$ seeds for each dataset setting.
Results We first compare how IRLEED improves over IQ, with online access to environment interactions. We provide these results in Tables 1 and 2, which show the return of the recovered policy relative to the return of the best demonstrator in that dataset. We can see that IRLEED outperforms its ablated counterpart, IQ, under all $3$ suboptimal dataset settings (not including the clean dataset), for both environments. Furthermore, we see that in the more complex Lunar Lander environment, IQ can not perform as well as the best demonstrator, whereas IRLEED consistently performs on par or better than the best demonstrator.
To have a fair comparison between IRLEED and its behaviour cloning counterpart, ILEED, we repeat these experiments without online access to environment interactions. Firstly, we can see that from Tables 1 and 2 that offline IRLEED outperforms ILEED, under all $3$ suboptimal dataset settings, for both environments. We can see that although ILEED has comparable performance in the simpler Cartpole environment, it performs poorly in the more complex Lunar Lander environment. Moreover, we can see from Table 2 that ILEED shows diminished performance when demonstrators have low accuracy $\epsilon$, signifying that it can not account for the reward biases present in these settings.
Table 1: Cartpole
We list the average return of the recovered policies (IQ, IRLEED, ILEED), relative to the return of the best demonstrator. This is provided for $4$ dataset settings (Prec., Acc.), where we list the mean performance of all demonstrators relative to the best demonstrator (Mean). The last row (H,H) corresponds to a clean dataset. The return of the best demonstrator corresponding to each setting (top to bottom) is: $260,500,411,500$. Setting Online Offline Prec. Acc. Mean IQ IRLEED ILEED IQ IRLEED L L $0.58$ $0.97$ 1.55 $1.38$ $0.96$ 1.56 L H $0.70$ $0.92$ 1.00 $1.00$ $0.96$ 1.00 H L $0.64$ $1.03$ 1.09 $1.05$ $0.95$ 1.05 H H $1.00$ $0.99$ $1.00$ $0.92$ $1.00$ $1.00$
Table 2: Lunar Lander
See Table 1. The return of the best demonstrator corresponding to each setting (top to bottom) is: $171,213,166,277$. Setting Online Offline Prec. Acc. Mean IQ IRLEED ILEED IQ IRLEED L L $0.46$ $0.68$ 1.28 $0.84$ $0.81$ 1.34 L H $0.54$ $0.84$ 1.14 $1.06$ $1.05$ 1.18 H L $0.57$ $0.94$ 1.30 $0.85$ $0.99$ 1.14 H H $0.86$ $0.93$ 0.98 $1.01$ $0.99$ $0.98$
5.3 Atari with Human Data
For our final experiment, we used two Atari environments, Space Invaders, and Qbert, utilizing both simulated and human demonstrations. Using these two environments, we compared the performance of IRLEED, IQ, and ILEED in the offline setting.
Setup Our demonstrations were comprised of two unique sources, dataset A: simulated data using a pretrained expert policy and dataset B: collected data using adept human players (Kurin et al., 2017). Since the human dataset B was acquired using screenshots of web-browsers, there is a small mismatch in sampling frequencies and colors when compared to the simulated environment used in dataset A. This provides a unique challenge akin to crowd-sourcing data from varying sources, where the goal is to see if we can achieve improved performance when training on the combined dataset AB as opposed to dataset A alone. Note that we followed the same steps outlined in the previous Section 5.2 to implement the three methods and setup the training and evaluation procedures. The results shown are averaged over $5$ seeds.
Results With this setup, we would like to test if IRLEED can improve the crowd-sourcing capability of IQ, and compare its performance to ILEED. We provide these results for both environments in Table 3, which lists the return of the recovered policies. Notably, all methods do poorly when trained exclusively on dataset B, due to the mismatch between the simulation environment and the data. On the other hand, utilizing dataset A is enough to reach decent performance. However, we can see that IRLEED is the only method that significantly improves its performance when utilizing the combined dataset AB as opposed to dataset A.
Table 3: Utilizing Demonstrations from Varying Sources
We list the average return of the recovered policies (IQ, IRLEED, ILEED). This is provided for $3$ dataset settings, simulated dataset A, human dataset B (Kurin et al., 2017), and their combination AB. Overall we see IRLEED shows a $16\%$ improvement over IQ, and a $41\%$ improvement over ILEED in the combined AB setting.
Dataset Space Invaders Qbert
ILEED IQ IRLEED ILEED IQ IRLEED
AB $802$ $755$ 910 $5511$ $8248$ 9336
A $783$ $754$ $768$ $5926$ $8092$ $8049$
B $110$ $179$ $123$ 00 $809$ $1087$
6 Discussion
In summary, this paper addresses the challenges of leveraging suboptimal and heterogeneous demonstrations in IL by introducing IRLEED, a novel IRL framework. By integrating a general model of demonstrator suboptimality that accounts for reward bias and action variance, with a maximum entropy IRL framework, IRLEED successfully mitigates the limitations inherent in standard IL algorithms, which often fail to account for the diversity and imperfections in real-world data. We empirically demonstrate that IRLEED can be used alongside existing IRL techniques to improve their performance in the presence of suboptimal and heterogeneous demonstrations. Moreover, we show that IRLEED outperforms its behavior cloning counterpart ILEED in the offline setting, demonstrating the importance of utilizing a dynamics-aware approach that considers reward bias. Although we have shown that IRLEED improves over IRL in a multitude of settings, we have not provided guarantees that state whether IRLEED will recover a better reward estimate compared to IRL. One direction for future work could be to provide more theoretical analysis comparing IRLEED to IRL under simpler settings with strict assumptions.
7 Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Implementation Details
@@
Below we detail the experimental setups utilized in Section 5, and provide information on the computational resources used for our work. Note that we provide a working implementation of the Gridworld experiments found in Section 5.1 as part of our supplementary material, and plan to release a public implementation of IRLEED with all experiments and datasets used after publication.
A.1 Gridworld
@@
To implement both IRLEED and IRL for the Gridworld experiments in Section 5.1, we utilized stochastic value iteration on the reward estimate $\theta$ to infer the policy $\pi$, and Monte Carlo simulations of the MDP to estimate the feature expectations required to optimize $\theta$, repeating both steps until convergence. As stated, when using IRLEED, we only updated $\epsilon$ and $\beta$ after $\theta$ converged, and repeated this for two iterations. For stochastic value iteration, we used a discount factor of $0.9$, a maximum horizon of $100$ timesteps, and a convergence criteria of $41e-$ 4 on the inferred state action values. For Monte Carlo simulations, we used $100$ episode samples. For optimization, we used stochastic gradient ascent with learning rates of $0.05,0.1,0.2$ for $\beta$, $\epsilon$, $\theta$, respectively, and a convergence criteria of $41e-$ 4 on $\theta$. For IRL, we initialized $\theta=[0.1]^{k}$, and set $\beta_{i}=1$ and $\epsilon_{i}=[0]^{k}$ to constants for all demonstrators, removing their effect. For evaluation, we used the policy recovered from the final estimate of $\theta$, sampling $100$ episodes to measure the mean reward.
@@
A.2 Simulated Control Tasks
To implement IQ and setup the training and evaluation procedures for the simulated control tasks in Section 5.2, we utilized the codebase provided by the authors. [^2] https://github.com/Div99/IQ-Learn. For all experiments, we utilized the provided hyperparameters for IQ. To implement ILEED, we followed the codebase provided by the authors, implementing the modified behaviour cloning loss alongside the IQ implementation. [^3] https://github.com/Stanford-ILIAD/ILEED For IRLEED, we built on top of the aforementioned IQ implementation by adding a learnable parameter $\beta_{i}$, and an additional critic network $\epsilon_{i}$, for each demonstrator $i$. This way, we sampled suboptimal policies according to: $\pi_{i}(a|s)\propto\exp(\beta_{i}(Q(s,a)+\alpha\epsilon_{i}(s,a))$. All three methods utilized the same architecture for the critic networks, and followed the original hyperparamters used for the IQ implementation. For ILEED, the critic network was used to compute action probabilities using a final softmax layer. For IRLEED and ILEED, $\beta_{i}$ was initialized to the same inverse temperature parameter used in the IQ implementation, and the learning rate was set to $45e-$ 4. For IRLEED, the additional critic networks $\epsilon_{i}(s,a)$ were initialized with the same procedure as the critic network $Q(s,a)$, utilized the same learning rate divided by $10$, and utilized the scaling factor $\alpha=0.01$. We trained all methods using $50,000$ and $100,000$ time steps for Cartpole and Lunar Lander, respectively. For evaluation, we utilized the final policy, sampling $300$ episodes to measure the mean reward.
@@ A.2 Simulated Control Tasks
A.3 Atari with Human Data
As aforementioned, the implementation for the Atari experiments in Section 5.3 was identical to the one described above in Section A.2. For training, we utilized $1,000,000$ time steps for all methods. For IRLEED, we did not update the error critic networks $\epsilon_{i}(s,a)$ or the temperature parameters $\beta_{i}$ for the first $300,000$ timesteps. For evaluation, we sampled $10$ episodes from the policies recovered during the last $5$ epochs of training, where each epoch corresponded to $5,000$ time steps.
The two datasets used are publicly available, where dataset A: simulated data using a pretrained expert policy, was provided alongside the original IQ implementation (Garg et al., 2021) and dataset B: collected data using adept human players, was part of a larger human study (Kurin et al., 2017). For Space Invaders, dataset A contained $5$ trajectories with an average return of $1285$, whereas dataset B contained $5$ trajectories with an average return of $1801$. For Qbert, dataset A contained $10$ trajectories with an average return of $14760$, whereas dataset B contained $10$ trajectories with an average return of $17340$. We ran this experiments using the three combinations of these datasets, averaging our results over $5$ seeds.
A.4 Computational Resources
All of the experiments were performed on an internal cluster containing the 20C/40T Intel Xeon Silver 4114 CPU, 64GB RAM, and 4×GTX-1080 GPUs. Obtaining results for the Gridworld experiment with $121$ datasets, each with $100$ seeds, took roughly $3$ days. Obtaining results for the simulated control tasks, Cartpole and Lunar Lander, took roughly $4$ days each. Training models on the Atari environments took $40$ hours per seed for IQ and IRLEED, and $20$ hours per seed for ILEED.

### v2

@@ Appendix A Extension to Generalized IRL
where $\psi$ is a convex reward regularizer that smoothly penalizes differences between occupancy measures to prevent overfitting (Ho and Ermon 2016). A naive solution to this nested min-max objective involves an outer loop learning the reward, and an inner loop executing maximum entropy RL with this reward to find the optimal policy. As before, the optimal policy under reward $r$ takes the form of the soft Bellman policy defined in Equation 3, where $r(s,a)$ replaces the linear reward $\theta^{\top}f(s,a)$. The description of IRLEED in the generalized setting follows directly.
@@
where in practice, the unknown functions $r$ and $\epsilon=\{\epsilon_{i}\}_{i=1}^{N}$ can be parameterized by neural networks. Note that under a constant reward regularizer $\psi$, if we express the rewards as $r_{i}(s,a)=(\theta+\epsilon_{i})^{\top}f(s,a)$, then the gradients of the above objective are identical to the ones described in Lemma 4.2. We note that in addition to the reward regularization used in Eq. (14), we use $\ell_{2}$-regularization on the output of neural networks $\{\epsilon_{i}\}_{i=1}^{N}$ to model the assumption that each demonstrator $i$ has a reward $r_{i}$ that varies slightly from the common reward $r$. One direction for future analysis would be to experiment with other regularization techniques, and see which assumptions are better for different kinds of data.
Appendix B Implementation Details
@@
Below we detail the experimental setups utilized in Section 5, and provide information on the computational resources used for our work. Note that we provide the implementation of IRLEED used during the Gridworld experiments, while referencing implementations used for other experiments below.
Gridworld
@@
To implement both IRLEED and IRL for the Gridworld experiments, we utilized stochastic value iteration on the reward estimate $\theta$ to infer the policy $\pi$, and Monte Carlo simulations of the MDP to estimate the feature expectations required to optimize $\theta$, repeating both steps until convergence. As stated, when using IRLEED, we only updated $\epsilon$ and $\beta$ after $\theta$ converged, and repeated this for two iterations. For stochastic value iteration, we used a discount factor of $0.9$, a maximum horizon of $100$ timesteps, and a convergence criteria of $41e-$ 4 on the inferred state action values. For Monte Carlo simulations, we used $100$ episode samples. For optimization, we used stochastic gradient ascent with learning rates of $0.05,0.1,0.2$ for $\beta$, $\epsilon$, $\theta$, respectively, and a convergence criteria of $41e-$ 4 on $\theta$. For IRL, we initialized $\theta=[0.1]^{k}$, and set $\beta_{i}=1$ and $\epsilon_{i}=[0]^{k}$ to constants for all demonstrators, removing their effect. For evaluation, we used the policy recovered from the final estimate of $\theta$, sampling $100$ episodes to measure the mean reward.
@@
Simulated Control Tasks
To implement IQ and setup the training and evaluation procedures for the simulated control tasks, we utilized the codebase provided by the authors. [^2] https://github.com/Div99/IQ-Learn For all experiments, we utilized the provided hyperparameters for IQ. To implement ILEED, we followed the codebase provided by the authors, implementing the modified behavior cloning loss alongside the IQ implementation. [^3] https://github.com/Stanford-ILIAD/ILEED For IRLEED, we built on top of the aforementioned IQ implementation by adding a learnable parameter $\beta_{i}$, and an additional critic network $\epsilon_{i}$, for each demonstrator $i$. This way, we sampled suboptimal policies according to: $\pi_{i}(a|s)\propto\exp(\beta_{i}(Q(s,a)+\alpha\epsilon_{i}(s,a))$. All three methods utilized the same architecture for the critic networks, and followed the original hyperparamters used for the IQ implementation. For ILEED, the critic network was used to compute action probabilities using a final softmax layer. For IRLEED and ILEED, $\beta_{i}$ was initialized to the same inverse temperature parameter used in the IQ implementation, and the learning rate was set to $45e-$ 4. For IRLEED, the additional critic networks $\epsilon_{i}(s,a)$ were initialized with the same procedure as the critic network $Q(s,a)$, utilized the same learning rate divided by $10$, a scaling factor of $\alpha=0.01$, and an $\ell_{2}$-regularization weight of $21e-$ 2. We trained all methods using $50,000$ and $100,000$ time steps for Cartpole and Lunar Lander, respectively. For evaluation, we utilized the final policy, sampling $300$ episodes to measure the mean reward.
@@
Atari with Human Data
As aforementioned, the implementation for the Atari experiments was identical to the one described above in Appendix B. For training, we utilized $1,000,000$ time steps for all methods. For IRLEED, we did not update the error critic networks $\epsilon_{i}(s,a)$ or the temperature parameters $\beta_{i}$ for the first $300,000$ timesteps. For evaluation, we sampled $10$ episodes from the policies recovered during the last $5$ epochs of training, where each epoch corresponded to $5,000$ time steps.
The two datasets used are publicly available, where dataset A: simulated data using a pretrained expert policy, was provided alongside the original IQ implementation (Garg et al. 2021) and dataset B: collected data using adept human players, was part of a larger human study (Kurin et al. 2017). For Space Invaders, dataset A contained $5$ trajectories with an average return of $1285$, whereas dataset B contained $5$ trajectories with an average return of $1801$. For Qbert, dataset A contained $10$ trajectories with an average return of $14760$, whereas dataset B contained $10$ trajectories with an average return of $17340$. We ran this experiments using the three combinations of these datasets, averaging our results over $5$ seeds.
Mujoco
To train on the continuous control Mujoco environments, we used the SAC implementation of IQ provided by the authors. For IRLEED, we built on top of the aforementioned IQ implementation by adding a learnable parameter $\beta_{i}$, and appending an additional network $\epsilon_{i}$ to the actor for each demonstrator $i$, keeping the critic network identical. All three methods utilized the same architectures, and followed the original hyperparamters used for the IQ implementation. For IRLEED, $\beta_{i}$ was initialized to the same inverse temperature parameter used in the IQ implementation, and the learning rate was set to $51e-$ 5. The additional actor networks $\epsilon_{i}(s,a)$ were initialized with the same procedure as the actor network $\pi(s,a)$, utilized the same learning rate divided by $10$, a scaling factor of $\alpha=0.01$, and an $\ell_{2}$-regularization weight of $21e-$ 2. We trained all methods using $500,000$ time steps. For evaluation, we sampled $10$ episodes from the policies recovered during training, abd averaged our result over $15$ initializations.
To create the different datasets consisting of A, B, and C, we utilized pretrained policies. Utilizing the demonstrations provided by the IQ codebase, we ran the IQ algorithm for $1,000,000$ timesteps, $50,000$ timesteps, and $1,000$ timesteps, creating three policies with varying optimality (corresponding to datasets A, B, and C, respectively). When creating the combined datasets, each policy was used to generate $1000$ state-action pairs, some containing more than one trajectory.
Computational Resources
All of the experiments were performed on an shared cluster containing the 20C/40T Intel Xeon Silver 4114 CPU, 64GB RAM, and 4×GTX-1080 GPUs.